# Plugging Attention into Power Grids: Towards Transparent Forecasting

### v1

@@ 4.2 Experimental Settings
A simple yet effective way to quantify uncertainty in neural networks is to train multiple models independently with different random initializations and aggregate their predictions [25]. This approach, known as Deep Ensembles, captures epistemic uncertainty – i.e. uncertainty stemming from limited knowledge or data – by leveraging the variability in predictions induced by different parameter configurations. It is particularly well- suited when models exhibit high variance, than can happen if small amounts of data are available or the variables are highly correlated. Given that each forecasting model carries distinct strengths and inductive biases, we further exploit their complementarity through expert aggregation. In particular, we employ the ML-Poly algorithm [13], a robust online learning method that dynamically adjusts expert weights based on past forecasting performance. In our setting, aggregation is performed every $48$ time steps (i.e. once per daily prediction horizon), with weights updated daily to reflect recent accuracy. Formally, for time step $t$, let $x_{j,t}$ denote the forecast of expert $j$ and $p_{j,t}$ its weight; the aggregated prediction is given by: $\widehat{y}_{t}=\sum_{j=1}^{K}p_{j,t}x_{j,t}$ where $K$ is the number of experts. To avoid overfitting and reduce the complexity of model selection, we also consider uniform aggregation, which consists of a simple average over all expert forecasts. Despite its simplicity, this baseline often performs competitively – sometimes even outperforming the best individual expert – thanks to the regularizing effect of ensemble averaging which reduces the variance of predictions.
@@ 4.3 Results
On the French dataset, which has a coarser spatial resolution and includes more aggregated load profiles at the regional level, performance across GNN architectures is relatively homogeneous. With uniform aggregation (over 10 random initializations for each method), SAGE delivers the best RMSE ($873$ MW) and a competitive MAPE ($1.12\%$), while the vanilla GAT achieves the lowest MAPE ($1.11\%$), confirming its ability to balance expressiveness and stability. Simpler models such as GCN and TAG remain highly competitive, and the gains from more complex architectures like GATv2 and TransformerConv are marginal or even negative. These findings suggest that in moderately sized graphs with stable and aggregated targets, moderately expressive models can match or outperform more elaborate architectures, particularly when using uniform aggregation, which helps mitigate overfitting risks and reduces variance across experts. Baseline methods such as FF or SARIMA are clearly outperformed by GNNs, reaffirming the utility of leveraging graph structure in this forecasting context.
@@
In contrast, the UK dataset is characterized by a much finer granularity and a greater number of disaggregated nodes (residential substations), where learning is more sensitive to data sparsity and noise. Under these conditions, APPNP with uniform aggregation achieves the best results, reaching the lowest MAPE ($6.54\%$) and RMSE ($12.61$ MW), followed by GAT ($7.79\%$, $14.13$ MW) and SAGE ($8.19\%$, $14.87$ MW). In this data-scarce regime, complex attention-based architectures such as GATv2 and TransformerConv perform worse, likely due to their higher parameterization and the resulting overfitting. The strong performance of APPNP can be attributed to its propagation scheme with teleportation, which regularizes diffusion and stabilizes learning. GAT’s strong results confirm that attention mechanisms can still offer competitive accuracy, especially when combined with uniform averaging, which smooths predictions across models and limits overfitting. These observations reinforce the value of simple and stable GNN designs in fine-grained and noisy environments.
@@
Overall, the results highlight the importance of architectural simplicity and strong inductive biases in low-data, high-resolution scenarios, as well as the value of graph-aware models over global baselines. Across both datasets, the choice of aggregation strategy significantly influences performance, often surpassing the differences between GNN architectures themselves. On the French dataset, the Bottom Aggregation with learned weights consistently yields the best results, achieving the lowest MAPE ($1.01\%$) and RMSE ($789$ MW). This suggests that learning to combine expert predictions at the local (node) level before aggregating them captures more informative patterns than naive averaging. Similarly, on the UK dataset, where signals are more heterogeneous and data is sparser, Bottom Aggregation (learned) also outperforms other strategies with the lowest MAPE, with $7.68\%$, and RMSE, with $14.35$ MW. This supports the idea that local aggregation provides a finer control over uncertainty and variation at the node level before summarizing across the network. In contrast, Top Aggregation – which combines predictions after summation – tends to perform worse, especially in the UK setting, likely because it neglects node-level diversity. Moreover, while uniform weighting remains competitive, particularly in simpler models, learned aggregation offers systematic gains, underlining the benefit of adaptive combination rules in ensemble-based forecasting.
@@
Table 2: Average and best test performance per model – ’Unif.’ indicates uniform averaging of expert predictions, ’Agg.’ refers to online aggregation using the MLPol algorithm. ’Bottom’ aggregation aggregates at the node level before summation, while ’Top’ aggregation first sums expert predictions and then aggregates. ’*’ means the models were only ran once as as their training procedure is deterministic .
@@
Agg. 1.09 $839$ 8.51 14.91
SAGE $\mathbf{1.30\pm 0.09}$ $\underline{1.14}$ $\mathbf{998\pm 72}$ $\mathbf{862}$ $\underline{8.63\pm 0.84}$ $6.95$ $\underline{15.61\pm 1.49}$ $13.10$
Unif. 1.12 873 8.19 14.87
Agg. 1.14 $898$ 8.07 14.85
GAT $1.42\pm 0.23$ $\mathbf{1.13}$ $1052\pm 136$ $\underline{888}$ $10.01\pm 1.80$ $\underline{6.59}$ $18.11\pm 2.96$ $\underline{12.79}$
Unif. 1.11 883 7.79 14.13
Agg. 1.08 $871$ 8.29 14.72
@@
Cheby $\underline{1.39\pm 0.14}$ $1.23$ $\underline{1036\pm 91}$ $925$ $\underline{9.49\pm 1.12}$ $7.95$ $18.38\pm 1.83$ $15.66$
Unif. 1.16 878 8.60 17.23
@@
APPNP $1.48\pm 0.17$ $1.23$ $1091\pm 114$ $931$ $\mathbf{8.31\pm 1.53}$ $\mathbf{6.01}$ $\mathbf{15.41\pm 2.73}$ $\mathbf{11.01}$
Unif. 1.22 944 6.54 12.61
Agg. 1.20 $943$ 7.60 14.13
FF $3.53\pm 0.09$ $3.34$ $2501\pm 47$ $2411$ $10.71\pm 0.40$ 9.83 $18.48\pm 609$ 17.06
SARIMA $3.29^{*}$ $3.29$ $2653^{*}$ $2653$ $8.06^{*}$ $8.06$ $16.60^{*}$ $16.60$
Persistence–$1$ $5.79^{*}$ $5.79$ $4512^{*}$ $4512$ $8.86^{*}$ $8.86$ $17.20^{*}$ $17.20$
Persistence–$7$ $6.03^{*}$ $6.03$ $4656^{*}$ $4656$ $10.19^{*}$ $10.19$ $19.19^{*}$ $19.19$
Bottom Uniform 1.06 836 8.10 14.74
Bottom Aggregation 1.01 789 7.68 14.35
Top Uniform 1.08 844 8.09 14.74
@@
To gain interpretability from attention-based architectures, we analyze the attention weights learned by GATs. For each time step in the dataset, the model computes a full attention matrix over all pairs of regions, quantifying the relative importance of inter-nodal interactions for the prediction task. These matrices, dynamically updated across time, encode rich spatio-temporal patterns. For instance, two regions may exhibit similar attention profiles in spring and autumn due to analogous consumption behaviors, yet diverge significantly in summer or winter – particularly in regions affected by seasonal tourism. To explore the structure embedded in these matrices, we apply dimensionality reduction techniques such as Principal Component Analysis (PCA) and Uniform Manifold Approximation and Projection (UMAP). These methods are especially useful for uncovering latent patterns in high-dimensional data: PCA captures directions of maximal variance, offering a linear projection, while UMAP preserves local and global topological structures, revealing non-linear relationships. As illustrated in Fig. 3, both projections reveal a clear seasonal organization of the attention matrices. The color gradient, indicating the day of the year, reveals well-separated clusters corresponding to different periods (e.g. winter, summer). This seasonal structuring suggests that the attention mechanism implicitly captures underlying physical variables such as temperature, daylight variation, or regional activity cycles – providing an interpretable signal beyond raw consumption values. These findings reinforce the potential of attention-based models not only as powerful predictors, but also as tools for understanding spatio-temporal dynamics in energy consumption.
@@
(a) Explicit coupling: Layer $2$.
(b) Implicit coupling: Layer 00.
(c) Implicit coupling: Layer $1$.
(d) Implicit coupling: Layer $2$.
Figure 4: PCA of attention matrices on the synthetic dataset with two coupling modes ($\mathbf{A}_{1}$ or $\mathbf{A}_{2}$). Colors reflect the coupling matrix in effect. (a) Explicit coupling information: the mode is explicitly given to the model as a binary input. (b-d) Implicit coupling information: the mode in effect is indicated implicitly by the combined values of two input binary variables. The subfigures show the PCA of attention matrices across GAT layers.
@@
To evaluate the model’s ability to recover hidden structure from ambiguous input, we design a third scenario with ambiguous switching. In this case, the same switching pattern between $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ is used as before, but the information about which of those matrices is in effect each time is now implicitly encodes by two categorical variables. This setup introduces ambiguity by suggesting four categories where only two actual regimes exist. We perform a PCA of the learned attention vectors at each GAT layer. As shown in Fig. 4, the first layer exhibits no clear separation. In the second layer, three intermediate clusters emerge, corresponding to the four ambiguous exogenous categories. In the final layer, these are merged into two well-separated clusters, correctly matching the underlying coupling structure. This progression suggests that GATs can disentangle latent generative factors through hierarchical attention refinement, even in the presence of misleading categorical inputs.
@@ 5 Conclusion
This paper investigates the use of attention-based graph neural networks for electricity consumption forecasting, with a strong focus on interpretability. We show that attention mechanisms not only improve forecasting accuracy on certain datasets, but also provide insights into time-varying regional dependencies. Our analysis of attention weight distributions reveals interpretable seasonal and meteorological patterns, which are not captured by classical black-box models. Interestingly, our results also show that simpler architectures such as GCN and APPNP often outperform deeper or more complex models (e.g. GATv2, TransformerConv), especially in highly disaggregated, low-data settings such as the UK residential network. These findings emphasize the importance of model simplicity and strong inductive biases in real-world forecasting scenarios. Finally, we show that ensemble techniques – particularly bottom-up aggregations with learned or uniform weights – offer a practical and effective strategy to improve robustness.
Future work may include learning graph structure jointly with forecasts and further exploring causal interpretations of attention mechanisms.

### v2

@@ 4.2 Experimental Settings
A simple yet effective way to quantify uncertainty in neural networks is to train multiple models independently with different random initializations and aggregate their predictions [26]. This approach, known as Deep Ensembles, captures epistemic uncertainty —i.e. uncertainty stemming from limited knowledge or data— by leveraging the variability in predictions induced by different parameter configurations. It is particularly well-suited when models exhibit high variance, than can happen if small amounts of data are available or the variables are highly correlated. Given that each forecasting model carries distinct strengths and inductive biases, we further exploit their complementarity through expert aggregation. In particular, we employ the ML-Poly algorithm [14], a robust online learning method that dynamically adjusts expert weights based on past forecasting performance. In our setting, aggregation is performed every $48$ time steps (i.e. once per daily prediction horizon), with weights updated daily to reflect recent accuracy. Formally, for time step $t$, let $x_{j,t}$ denote the forecast of expert $j$ and $p_{j,t}$ its weight; the aggregated prediction is given by: $\widehat{y}_{t}=\sum_{j=1}^{K}p_{j,t}x_{j,t}$ where $K$ is the number of experts. To avoid overfitting and reduce the complexity of model selection, we also consider uniform aggregation, which consists of a simple average over all expert forecasts. Despite its simplicity, this baseline often performs competitively —sometimes even outperforming the best individual expert— thanks to the regularizing effect of ensemble averaging which reduces the variance of predictions.
@@ 4.3 Results
On the French dataset, which has a coarser spatial resolution and includes more aggregated load profiles at the regional level, performance across GNN architectures is relatively homogeneous. With uniform aggregation (over 10 random initializations for each method), SAGE delivers the best RMSE ($873$ MW) and a competitive MAPE ($1.12\%$), while the vanilla GAT achieves the lowest MAPE ($1.11\%$), confirming its ability to balance expressiveness and stability. Simpler models such as GCN and TAG remain highly competitive, and the gains from more complex architectures like GATv2 and TransformerConv are marginal or even negative. These findings suggest that in moderately sized graphs with stable and aggregated targets, moderately expressive models can match or outperform more elaborate architectures, particularly when using uniform aggregation, which helps mitigate overfitting risks and reduces variance across experts. Baseline methods such as FF, SARIMA, or persistence approaches are consistently outperformed by GNN-based models across both datasets, particularly in terms of RMSE. This highlights the benefit of exploiting graph structure for load forecasting, even though simpler baselines remain competitive in some cases (e.g., FF on the French dataset, and SARIMA on the UK dataset).
@@
In contrast, the UK dataset is characterized by a much finer granularity and a greater number of disaggregated nodes (residential substations), where learning is more sensitive to data sparsity and noise. Under these conditions, APPNP with uniform aggregation achieves the best results, reaching the lowest MAPE ($6.54\%$) and RMSE ($12.61$ MW), followed by GAT ($7.79\%$, $14.13$ MW) and SAGE ($8.19\%$, $14.87$ MW). In this data-scarce regime, complex attention-based architectures such as GATv2 and TransformerConv perform worse, likely due to their higher parameterization and the resulting overfitting. The strong performance of APPNP can be attributed to its propagation scheme with teleportation, which regularizes diffusion and stabilizes learning. GATs’ strong results confirm that attention mechanisms can still offer competitive accuracy, especially when combined with uniform averaging, which smooths predictions across models and limits overfitting. These observations reinforce the value of simple and stable GNN designs in fine-grained and noisy environments.
@@
Overall, the results highlight the importance of architectural simplicity and strong inductive biases in low-data, high-resolution scenarios, as well as the value of graph-aware models over global baselines. Across both datasets, the choice of aggregation strategy significantly influences performance, often surpassing the differences between GNN architectures themselves. On the French dataset, the bottom aggregation consistently yields the best results, achieving the lowest MAPE ($1.01\%$) and RMSE ($789$ MW). This suggests that learning to combine expert predictions at the local (node) level before aggregating them captures more informative patterns than naive averaging. Similarly, on the UK dataset, where signals are more heterogeneous and data is sparser, bottom aggregation also outperforms other strategies with the lowest MAPE, with $7.68\%$, and RMSE, with $14.35$ MW. This supports the idea that local aggregation provides a finer control over uncertainty and variation at the node level before summarizing across the network. In contrast, top aggregation —which combines predictions after summation— tends to perform worse, especially in the UK setting, likely because it neglects node-level diversity. Moreover, while uniform weighting remains competitive, particularly in simpler models, learned aggregation offers systematic gains, underlining the benefit of adaptive combination rules in ensemble-based forecasting.
@@
Table 2: Average and best test performance per model – ‘Unif.’ indicates uniform averaging of expert predictions, ‘Agg.’ refers to online aggregation using the MLPol algorithm. ‘Bottom’ aggregation aggregates at the node level before summation, while ‘Top’ aggregation first sums expert predictions and then aggregates. ‘*’ means the models were only ran once as as their training procedure is deterministic.
@@
Agg. $\underline{1.09}$ $\mathbf{839}$ $8.51$ $14.91$
SAGE $30\pm 0.09 1.1$41.14 $998\pm 72$ $862$ $8.63\pm 0.84$ $6.95$ $15.61\pm 1.49$ $13.10$
Unif. $\underline{1.12}$ $\mathbf{873}$ $8.19$ $14.87$
Agg. $1.14$ $898$ $\underline{8.07}$ $14.85$
GAT $42\pm 0.23 1.1$31.13 $1052\pm 136$ $888$ $10.01\pm 1.80$ $6.59$ $18.11\pm 2.96$ $12.79$
Unif. $\mathbf{1.11}$ $883$ $\underline{7.79}$ $\underline{14.13}$
Agg. $\mathbf{1.08}$ $\underline{871}$ $8.29$ $\underline{14.72}$
@@
Cheby $1.39\pm 0.14$ $1.23$ $1036\pm 91$ $925$ $9.49\pm 1.12$ $7.95$ $18.38\pm 1.83$ $15.66$
Unif. $1.16$ $\underline{878}$ $8.60$ $17.23$
@@
APPNP $1.48\pm 0.17$ $1.23$ $1091\pm 114$ $931$ $8.31\pm 1.53$ $6.01$ $5.41\pm 2.73 11$.0111.01
Unif. $1.22$ $944$ $\mathbf{6.54}$ $\mathbf{12.61}$
Agg. $1.20$ $943$ $\mathbf{7.60}$ $\mathbf{14.13}$
$F$ $1.50\pm 0.06$ $1.41$ $1101\pm 31$ $1044$ $10.71\pm 0.40$ $9.83$ $18.48\pm 609$ $17.06$
SARIMA $^{*} 3.293.29$ $^{*} 26532653$ $^{*} 8.068.06$ $^{*} 16.6016.60$
Persistence–11 $^{*} 5.795.79$ $^{*} 45124512$ $^{*} 8.868.86$ $^{*} 17.2017.20$
Persistence–77 $^{*} 6.036.03$ $^{*} 46564656$ $^{*} 10.1910.19$ $^{*} 19.1919.19$
Uniform $1.06$ $836$ $8.10$ $14.74$
Bottom Aggregation $\mathbf{1.01}$ $\mathbf{789}$ $\mathbf{7.68}$ $\mathbf{14.35}$
@@
To gain interpretability from attention-based architectures, we analyze the attention weights learned by GATs. For each time step in the dataset, the model computes a full attention matrix over all pairs of regions, quantifying the relative importance of inter-nodal interactions for the prediction task. These matrices, dynamically updated across time, encode rich spatio-temporal patterns. For instance, two regions may exhibit similar attention profiles in spring and autumn due to analogous consumption behaviors, yet diverge significantly in summer or winter —particularly in regions affected by seasonal tourism. To explore the structure embedded in these matrices, we apply dimensionality reduction techniques such as Principal Component Analysis (PCA) and Uniform Manifold Approximation and Projection (UMAP). These methods are especially useful for uncovering latent patterns in high-dimensional data: PCA captures directions of maximal variance, offering a linear projection, while UMAP preserves local and global topological structures, revealing non-linear relationships. As illustrated in Fig. 3, both projections reveal a clear seasonal organization of the attention matrices. The color gradient, indicating the day of the year, reveals well-separated clusters corresponding to different periods (e.g. winter, summer). This seasonal structuring suggests that the attention mechanism implicitly captures underlying physical variables such as temperature, daylight variation, or regional activity cycles —providing an interpretable signal beyond raw consumption values. These findings reinforce the potential of attention-based models not only as powerful predictors, but also as tools for understanding spatio-temporal dynamics in energy consumption.
@@
(a) Explicit coupling: Layer 22.
(b) Implicit coupling: Layer 0.
(c) Implicit coupling: Layer 11.
(d) Implicit coupling: Layer 22.
Figure 4: PCA of attention matrices on the synthetic dataset with two coupling modes ($\mathbf{A}_{1}$ or $\mathbf{A}_{2}$). Colors reflect the coupling matrix in effect. (a) Explicit coupling information: the mode is explicitly given to the model as a binary input. (b—d) Implicit coupling information: the mode in effect is indicated implicitly by the combined values of two input binary variables. The subfigures show the PCA of attention matrices across GAT layers.
@@
To evaluate the model’s ability to recover hidden structure from ambiguous input, we design a third scenario with ambiguous switching. In this case, the same switching pattern between $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$ is used as before, but the information about which of those matrices is in effect each time is now implicitly encodes by two categorical variables. This setup introduces ambiguity by suggesting four categories where only two actual regimes exist. We perform a PCA of the learned attention vectors at each GAT layer. As shown in Fig. 4, the first layer exhibits no clear separation. In the second layer, three intermediate clusters emerge, corresponding to the four ambiguous exogenous categories. In the final layer, these are merged into two well-separated clusters, correctly matching the underlying coupling structure. This progression suggests that GATs can disentangle latent generative factors through hierarchical attention refinement, even in the presence of misleading categorical inputs. This property is particularly valuable in real-world forecasting scenarios, where exogenous variables may be noisy, redundant, or poorly specified —as is often the case with categorical features in energy systems. In such cases, the model’s ability to infer underlying structure without explicit supervision can enhance robustness and generalization. Moreover, it opens promising avenues for model interpretability: analyzing how attention distributions evolve across layers may help identify meaningful patterns or changes of regime, even when external signals are ambiguous or incomplete. Ultimately, this illustrates how attention-based GNNs can support both prediction and explanation, by learning to filter out misleading information and emphasizing latent, structure-relevant features.‘
@@ 5 Conclusion
This paper investigates the use of attention-based graph neural networks for electricity consumption forecasting, with a strong focus on interpretability. We show that attention mechanisms not only improve forecasting accuracy on certain datasets, but also provide insights into time-varying regional dependencies. Our analysis of attention weight distributions demonstrates that GAT-based architectures can partially recover meaningful seasonal and meteorological patterns from the data, and highlights how attention mechanisms provide a structured and interpretable lens through which temporal and spatial dependencies can be explored. Interestingly, our results also show that simpler architectures such as GCN and APPNP often outperform deeper or more complex models (e.g. GATv2, TransformerConv), especially in highly disaggregated, low-data settings such as the UK residential network. These findings emphasize the importance of model simplicity and strong inductive biases in real-world forecasting scenarios. Finally, we show that ensemble techniques —particularly bottom-up aggregations— offer a practical and effective strategy to improve robustness.